OpenStudy (anonymous):

Seeking for the right answer..help me please.. The question is about differentiate with respect to x. Given, y=e with power 3x cosh x square

OpenStudy (unklerhaukus):

do you mean \[y = e^{3x}\cosh x^2\]or\[y = e^{3x}\cosh^2 x\]?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

is the square on the coshine or on the x ?

OpenStudy (anonymous):

on the x

OpenStudy (unklerhaukus):

ok so \[y = e^{3x}\cosh x^2\] and you have to take the derivative with respect to \(x\)

OpenStudy (unklerhaukus):

Do you remember the product rule of differentiation?

OpenStudy (anonymous):

is it differentiate both? so, dy/dx =e^3x and d/dx cosh x^2

OpenStudy (unklerhaukus):

The product rule states that if a function \(y(x)\) is composed of a product of functions \(f(x),\, g(x)\)\[y(x)=f\cdot g\] The derivative will be equal to the sum: \[y'= f'g+fg'\]

OpenStudy (unklerhaukus):

so let \(f(x) = e^{3x}\) and find the derivative \(f'= \quad. . .\) and let \(g(x)=\cosh(x^2)\) and find its derivative also \(g' = \quad...\)

OpenStudy (anonymous):

if we differentiate cosh x^2, is it get sinh x^2?

OpenStudy (anonymous):

and differentiate e^3x will remain same, e^3x?

OpenStudy (unklerhaukus):

dont forget the chain rule d/dx ( cosh (x^2) ) = sinh (x^2) * d/dx ( x^2 ) and also d/dx ( e^3x ) = e^3x * d/dx ( 3x )

OpenStudy (anonymous):

ohh i see..

OpenStudy (unklerhaukus):

\[y(x)=(f\circ g)(x)=f(g(x))\] when taking the derivative of a function of a function, you have to multiply the derivative of the outside function, by the derivative of the inside function \[y'=(f'\circ g)\cdot g'\]

OpenStudy (unklerhaukus):

for example \[y = e^{x^2}\\ y'=2x\cdot e^{x^2}\]

OpenStudy (anonymous):

i see.. if differentiate e^x will remain same .. is it right?

OpenStudy (unklerhaukus):

yes d/dx ( e^x ) = e^x * d/dx ( x ) = e^x * 1 = e^x

OpenStudy (anonymous):

i see.. so the answer for this question is 3e^3x sinh x^2 x^2

OpenStudy (anonymous):

hehe is it right? or can simply?

OpenStudy (anonymous):

ehh sorry, the last is 2x, not x^2

OpenStudy (unklerhaukus):

From the product formula\[y=f\cdot g\implies y'= f'g+fg'\] we let \(f=e^{3x}\) and \(g=\cosh(x^2)\) what did you get for \(f'\) and \(g'\) ?

OpenStudy (anonymous):

f' = 3e^3x and g' = 2x sinh x^2

OpenStudy (unklerhaukus):

that's right! now the derivative of the function \(y\) is \[y'= f'g+fg'\]

OpenStudy (anonymous):

yess.. need to plus both?

OpenStudy (unklerhaukus):

\[y'=3e^{3x}\cdot\cos(x^2)+e^{3x}\cdot2x\sinh(x^2)\]

OpenStudy (unklerhaukus):

you can simplify this if you factor it a little

OpenStudy (anonymous):

how..? i can't see :D

OpenStudy (unklerhaukus):

there is a common factor to both terms in the sum

OpenStudy (anonymous):

e^3x (3cos(x^2) + 2xsinh(x^2) ?

OpenStudy (anonymous):

x^2 can taken out?

OpenStudy (unklerhaukus):

That's it, your done

OpenStudy (unklerhaukus):

the x^2's are trapped the hyperbolic functions

OpenStudy (anonymous):

thats all? hee thanks a lot for help ^^

OpenStudy (anonymous):

i see